OpenStudy (anonymous):

You roll a blue die and a yellow die. What is P(the sum of the dice is at least 6 | the blue die shows a 4)?

OpenStudy (anonymous):

give me one minutes

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

two die = 12 12/6=2 12/4=3 so 2/3

OpenStudy (anonymous):

Thank you ! @fabo

OpenStudy (anonymous):

welcome

OpenStudy (mathmate):

S=the sum of the dice is at least 6 B= the blue die shows a 4 want to find: P(the sum of the dice is at least 6 | the blue die shows a 4) =P(S|B) \( P(S\cap B) =|{(2,4),(3,4),(4,4),(5,4),(6,4)}|\ /\ 36 \\ ={\large \frac{5}{36}} \\ P(B)=\frac{1}{6} \\ \ \\ But\ since \\ P(S|B)=\frac{P(S\cap B)}{P(B)} \\ = {\large \frac{\frac{5}{36}}{\frac{1}{6}} } \\ = \large \frac{5}{6} \)

OpenStudy (anonymous):

thanks! @mathmate

OpenStudy (mathmate):

You're welcome!